# Scaffold-SLAM: Structured 3D Gaussians for Simultaneous Localization and Photorealistic Mapping

### v1

@@ Abstract
![[Uncaptioned image]](extracted/6121471/fig/fig1/tum_rgbd/GS-ICP/01317.png)
![[Uncaptioned image]](extracted/6121471/fig/fig1/tum_rgbd/SplaTAM/01320.png)
![[Uncaptioned image]](extracted/6121471/fig/fig1/tum_rgbd/MonoGS/01317.png)
![[Uncaptioned image]](extracted/6121471/fig/fig1/tum_rgbd/Photo-SLAM/1341848025.834828.png)
![[Uncaptioned image]](extracted/6121471/fig/fig1/tum_rgbd/HQ-SLAM/30000_1296.jpg)
![[Uncaptioned image]](extracted/6121471/fig/fig1/tum_rgbd/GT/30000_1296.jpg)
![[Uncaptioned image]](extracted/6121471/fig/fig1/replica_off3_mono/Photo-SLAM/frame000100.jpg)
![[Uncaptioned image]](extracted/6121471/fig/fig1/replica_off3_mono/HQ-SLAM/30000_100.jpg)
![[Uncaptioned image]](extracted/6121471/fig/fig1/replica_off3_mono/GT/30000_100.jpg)
![[Uncaptioned image]](extracted/6121471/fig/fig1/euroc_stereo/MonoGS/01028.png)
![[Uncaptioned image]](extracted/6121471/fig/fig1/euroc_stereo/HQ-SLAM/30000_1827.jpg)

### v2

@@ Abstract
![[Uncaptioned image]](extracted/6234247/fig/fig1/tum_rgbd/GS-ICP/01317.png)
![[Uncaptioned image]](extracted/6234247/fig/fig1/tum_rgbd/SplaTAM/01320.png)
![[Uncaptioned image]](extracted/6234247/fig/fig1/tum_rgbd/MonoGS/01317.png)
![[Uncaptioned image]](extracted/6234247/fig/fig1/tum_rgbd/Photo-SLAM/1341848025.834828.png)
![[Uncaptioned image]](extracted/6234247/fig/fig1/tum_rgbd/HQ-SLAM/30000_1296.jpg)
![[Uncaptioned image]](extracted/6234247/fig/fig1/tum_rgbd/GT/30000_1296.jpg)
![[Uncaptioned image]](extracted/6234247/fig/fig1/replica_off3_mono/Photo-SLAM/frame000100.jpg)
![[Uncaptioned image]](extracted/6234247/fig/fig1/replica_off3_mono/HQ-SLAM/30000_100.jpg)
![[Uncaptioned image]](extracted/6234247/fig/fig1/replica_off3_mono/GT/30000_100.jpg)
![[Uncaptioned image]](extracted/6234247/fig/fig1/euroc_stereo/MonoGS/01028.png)
![[Uncaptioned image]](extracted/6234247/fig/fig1/euroc_stereo/HQ-SLAM/30000_1827.jpg)